Foreverangel:

What is the resistance of a parallel circuit with a 20-ohm resistor and a 10-ohm resistor?

AZ:

So do you know how to calculate resistance of a parallel circuit?

1 attachment
AZ:

If it was a series circuit, you would just add but for parallel it's different \( \dfrac{1}{R} = \dfrac{1}{R_1} + \dfrac{1}{R_2}\) so \( R_1\) is 20 and \( R_2\) is 10 can you solve for R?